find the zeros of the function h(x)=x^2+48x-100 by factoring
Answer by scott8148(6628)

h(x) = (x + 50)(x - 2)

0 = (x + 50)(x - 2)

0 = x + 50 ___ -50 = x

0 = x - 2 ___ 2 = x